### v1

# COUNTDOWN: Contextually Sparse Activation Filtering Out Unnecessary Weights in Down Projection

Abstract
The growing size of large language models has created significant computational inefficiencies. To address this challenge, sparse activation methods selectively deactivates non-essential parameters during inference, reducing computational costs in FFNN layers. While existing methods focus on non-linear gating mechanisms, we hypothesize that the sparsity of the FFNN layer lies globally in the form of a linear combination over its internal down projection matrix. Based on this insight, we propose two methods: M-CountDown, leveraging indirect coefficients, and D-CountDown, utilizing direct coefficients of the linear combination. Experimental results demonstrate that D-CountDown can omit 90% of computations with performance loss as low as 5.5% ideally, while M-CountDown provides a predictor-free solution with up to 29.4% better performance preservation compared to existing methods. Our specialized kernel implementations effectively realize these theoretical gains into substantial real-world acceleration.
COUNTDOWN: Contextually Sparse Activation Filtering Out
Unnecessary Weights in Down Projection
Jaewon Cheon Industrial and Management Engineering Korea University jaewon_cheon@korea.ac.kr Pilsung Kang* Industrial Engineering Seoul National University pilsung_kang@snu.ac.kr
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse applications, from handling specific tasks to orchestrating agent-based operations (OpenAI et al., 2024; DeepSeek-AI et al., 2024; Gemma Team et al., 2025). However, these advancements came at the cost of dramatically increased model sizes, creating enormous computational and resource demands.
The inference process has emerged as a particularly acute efficiency constraint, forming a critical bottleneck for deploying LLMs in practical applications. This inefficiency is further amplified by recent trends in test-time scaling, where models generate extensive reasoning, significantly increasing computational demands during inference (Jang et al., 2024; Deng et al., 2024). Consequently, research on LLM inference efficiency has intensified, aiming to reduce latency and memory consumption while preserving generation quality (Liu et al., 2024; Kwon et al., 2023; Cai et al., 2024).
In this context, sparse activation has emerged as a prominent strategy to improve FFNN layer efficiency in Transformer-based LLM (Liu et al., 2023; Lee et al., 2024; Akhauri et al., 2024; Alizadeh et al., 2024). Sparse activation methods dynamically identify and deactivate parameters unnecessary for a given input, thereby reducing computational load and accelerating inference. These methods are particularly beneficial since FFNN layers incur significant computational overhead in modern LLM architectures (Awasthi et al., 2024).
The zero-out gating property of ReLU (Agarap, 2019) creates extensive sparsity in FFNN layers by forcing a large portion of neurons to output zero (Mirzadeh et al., 2024). This natural sparsity makes computations associated with these zero-valued neurons completely redundant. Existing sparse activation methods leverage this property to identify and skip these unnecessary computations (Sun et al., 2024; Zhang et al., 2024). However, recent LLMs largely employ activations such as GeLU or SiLU (Hendrycks and Gimpel, 2016; Elfwing et al., 2018) with far less prevalent zero-out behavior (Mirzadeh et al., 2024), limiting these methods’ applicability. Further, Gated-MLP structures, now widely adopted as FFNN layers (Shazeer, 2020; Dauphin et al., 2017), introduce more complex parameter interactions than standard architectures. This invalidates the assumption that sparsity occurs only around non-linear activations.
To overcome these limitations, we propose an approach that defines sparsity from a global view, extending beyond the non-linear activations by reformulating the FFNN layer’s output as a weighted sum, as illustrated in Figure 1. Based on this approach, we derive two sparse activation methodologies: Mono-CountDown (M-CountDown) and Dual-CountDown (D-CountDown). M-CountDown identifies sparsity based on the output of a single weight matrix in Gated-MLP, while D-CountDown leverages two weight matrices. In evaluations, M-CountDown consistently outperforms the baseline method CATS (Lee et al., 2024), achieving up to 29.4% better performance preservation with comparable inference speed. D-CountDown attains greater efficiency gains, reducing computations by up to 90% in FFNN layers with performance loss as low as 5.5% under optimal conditions.
The contributions of this paper are as follows.
We introduce a novel theoretical framework that redefines sparsity through a weighted-sum perspective over down projection matrices, extending beyond the conventional focus on activation functions.
We demonstrate that analyzing coefficient vectors in the weighted sum enables superior sparsity decisions, resulting in two distinct approaches with complementary strengths.
We provide practical acceleration through optimized kernel implementations, enabling both methods to achieve substantial throughput improvements across multiple state-of-the-art LLM architectures.
2 Related Works
ReLU-based Sparse Activation
Early works on sparse activation primarily leveraged the property of ReLU to enhance computational efficiency. These approaches identified that ReLU activation functions naturally create substantial built-in sparsity by producing zeros for negative values (Li et al., 2023b). Several approaches have tried to detect these zero-valued activations to preemptively skip associated computations, as these neurons would have no impact on subsequent layers (Mirzadeh et al., 2024). Deja Vu (Liu et al., 2023) extended this concept by training lightweight predictors to anticipate which neurons would be zeroed out, further improving efficiency. While these methods showed impressive speed gains with minimal performance loss, their application faced significant constraints. Notably, these approaches were practical only on architectures explicitly designed with ReLU activations, limiting their applicability as LLMs increasingly adopted alternative activation functions (Akhauri et al., 2024).
Non-ReLU Sparse Activation
As LLM evolved to favor non-ReLU activation functions such as GeLU and SiLU, which rarely produce exact zeros, new methods emerged to extend sparsity benefits to these architectures. One direction involved ReLUfication techniques that replace non-ReLU functions with ReLU, enabling the reuse of existing sparsity strategies (Song et al., 2024b, 2025; Zhang et al., 2024; Alizadeh et al., 2024). Another approach, such as by CATS (Lee et al., 2024), redesigned sparsity criteria to identify and skip computations associated with near-zero activations rather than exact zeros. While these adaptations improved compatibility with modern LLM architectures, they remain fundamentally constrained by their narrow focus on local patterns around non-linear transformations, overlooking potential sparsity from a global perspective of the FFNN layer. This localized perspective may fail to fully capitalize on the potential sparsity distributed throughout modern Gated-MLP architectures, particularly considering the complex interactions among multiple weight matrices that define these structures.
3 Generalization of Sparse Activation
Problem Formulation
A Gated-MLP block consists of three weight matrices: . For this block, the input vector and the output vector are in . The computation involves intermediate states defined as in .
When no sparsification is applied, which we refer to as the Dense scenario, all parameters are activated, and the operation proceeds as follows:
where denotes a non-linear activation function, typically GeLU or SiLU.
We now introduce our sparsity propagation framework, establishing sparse activation from a global perspective. We can activate only a valuable subset of weight vectors, with a marginal performance loss. Specifically, sparse activation under our framework follows:
where denotes the column of indices of the weights selected for computation:
where is any function filtering effective .
Notably, when individual threshold functions are defined separately for each matrix, identical output can be achieved through the unified intersection :
Consequently, even when sparsifying just one matrix and keeping others dense, the computation remains equivalent to applying this unified across all matrices, which we denotes as shared-index property. Thus, if valuable sparsity patterns are identified in one matrix, they can propagate throughout the entire Gated-MLP.
A critical challenge, therefore, is defining the optimal filtering function to identify the most effective index set to preserve globally essential computations while significantly reduce computational overhead.
Limitation of Comparative Methodology
CATS (Lee et al., 2024) partially satisfies our sparsity propagation framework. It identifies sparsity by examining the activation magnitude , assuming activations squashed near zero indicate parameters to omit. Specifically, given a sparsity ratio , CATS computes a threshold via the operation, selecting a cutoff below which the lowest fraction of activations is excluded. Based on this threshold, CATS defines a sparse activation index as shown in Equation 6a.
CATS leverages the shared-index property. However, since the optimal might depend on factors beyond non-linear activation region, CATS is theoretically limited in propagating an optimal throughout the Gated-MLP. Additionally, although is large, if the corresponding is near zero, the final contributions become minimal, which ideally should be filtered out due to their elementwise product.
Threshold Variants
To overcome these limitations, we reformulate the Gated-MLP computation as a linear combination of the weight vectors, thereby exploring additional possibilities for defining as follows:
This reformulation allows us to interpret output as a weighted sum over row vectors, where coefficient reflects the -th row vector’s contribution to computation. The magnitude of these coefficients provides a natural metric for determining which parameters to activate, as they quantify each vector’s significance to the output.
Furthermore, since is calculated as the elementwise multiplication of and , these intermediate vectors can also serve as indirect coefficient signals. This generalized view reveals that each computation stage in the Gated-MLP can provide a distinct sparsity indicator, with selecting as the basis being equivalent to CATS’s approach.
Based on this view, we propose two variants of sparse activation that extend beyond prior approaches relying solely on the magnitude of . As shown in Equation 6, the first method, M-CountDown, applies thresholding directly to vector , while the second method, D-CountDown, applies thresholding to . For each method, thresholds and are calculated via and respectively.
These methods offer complementary strengths: M-CountDown provides practical implementation with minimal overhead by examining only one matrix multiplication, while D-CountDown can offer more precise sparsity identification through direct coefficients of the weighted sum. A detailed discussion of these methods follows in section 4
4 Realization of Sparse Activation
SPIdeal and SPPrac
In the previous section, we focused on establishing and the corresponding indicator that theoretically guarantee the safe omission of parameters. Ideally, if these indicators are tractable in real-time inference, we can achieve the upper-bound performance defined by the method. However, accessing the indicator and deriving from it is not trivial.
Given this constraint, we distinguish between two distinctive perspectives: SPIdeal examines the theoretical upperbound performance achievable by each method, assuming that filtering based on sparsity indicators incurs no computational overhead. SPPrac accounts for real-world deployment constraints, particularly the latency of identifying sparse activation patterns. It evaluates whether methods can deliver actual inference speedups when all practical overheads are considered.
The distinction is critical because methods with strong SPIdeal performance may not translate to SPPrac benefits if their practical implementation is computationally expensive. Conversely, focusing solely on SPPrac without understanding the theoretical SPIdeal limits can lead to suboptimal solutions that fail to approach the best possible performance. Effective sparse activation requires both identifying truly essential computations via SPIdeal and creating an efficient implementation to realize total computational savings through SPPrac.
Constructing SPPrac for CountDown
We now describe how to transform the theoretical SPIdeal formulations of M-CountDown and D-CountDown into efficient, practical SPPrac implementations.
For M-CountDown, the implementation is straightforward because its indicator depends only on the matrix . Therefore, its index set defined in its SPIdeal perspective can be obtained independently of other matrices in the Gated-MLP. This allows M-CountDown to operate without additional inference-time components, as computation over the remaining matrices can be selectively skipped based on .
To further reduce overhead, we avoid computing dynamically for each input. Instead, we approximate it with a layerwise constant estimated during a calibration phase with sampled inputs.
In contrast, implementing D-CountDown poses greater challenges because its indicator requires nearly the entire Gated-MLP computation, negating the advantages of sparse activation. To tackle this challenge, we train a lightweight predictor that estimates the optimal index set directly from input , avoiding the need to compute during inference. For each layer, the predictor outputs a score vector where:
Using this output, we define the predicted index set as and activate only the corresponding weight columns during inference.
For efficiency, the predictor must be highly accurate and computationally inexpensive during inference. Following (Liu et al., 2023; Alizadeh et al., 2024), we employ a low-rank approximator consisting of two matrices: and , minimizing computational overhead while preserving prediction accuracy. algorithm 1 details the complete training procedure.
Kernel Design
Once the sparse activation index set is determined, computation can be restricted to only the corresponding subset of weights, reducing the actual floating-point operation count (FLOPs). However, reducing FLOPs does not necessarily translate to improved inference latency. For instance, materializing an indexed weight matrix and performing standard vector-matrix multiplication may still reduce FLOPs, but at the cost of increased memory access (Song et al., 2024a; Xue et al., 2024). Therefore, sparse computation should avoid incurring excessive memory traffic solely for the sake of reducing arithmetic operations.
To address this, we implement custom kernels for both M-CountDown and D-CountDown using Triton (Tillet et al., 2019). The M-CountDown kernel builds upon CATS’s structure (Lee et al., 2024), but optimizes it by fusing the non-linear activation to reduce additional memory access. For D-CountDown, we design a kernel that efficiently supports predictor-based activation patterns. A naive implementation would require eight separate kernel launches for sparse computation: indexing and GEMV for each of the three matrices, plus non-linear activation and elementwise multiplication. Our implementation compresses this workload into just two kernels. This design ensures that FLOPs reductions directly translate into improved token throughput. Full implementation details and pseudocode are in algorithm 2 and algorithm 3.
InferenceMode | Llama-3.1-8B-Instruct | gemma-2-9b-it | Qwen2.5-14B-Instruct | phi-4 | ||||||||
Dense | ||||||||||||
Full | 0.616 | 0.645 | 0.674 | 0.655 | ||||||||
SPIdeal | ||||||||||||
DEJAVU | 0.314 | 0.315 | 0.322 | 0.360 | 0.360 | 0.360 | 0.379 | 0.382 | 0.385 | 0.398 | 0.405 | 0.396 |
CATS | 0.471 | 0.412 | 0.337 | 0.592 | 0.483 | 0.367 | 0.502 | 0.428 | 0.389 | 0.615 | 0.535 | 0.427 |
M-CountDown | 0.570 | 0.513 | 0.421 | 0.624 | 0.607 | 0.549 | 0.644 | 0.610 | 0.479 | 0.636 | 0.608 | 0.512 |
D-CountDown | 0.603 | 0.587 | 0.525 | 0.635 | 0.625 | 0.590 | 0.660 | 0.647 | 0.555 | 0.651 | 0.649 | 0.620 |
SPPrac | ||||||||||||
CATS | 0.504 | 0.450 | 0.350 | 0.605 | 0.502 | 0.360 | 0.556 | 0.478 | 0.390 | 0.633 | 0.591 | 0.448 |
M-CountDown | 0.571 | 0.528 | 0.447 | 0.632 | 0.617 | 0.588 | 0.651 | 0.624 | 0.535 | 0.639 | 0.620 | 0.555 |
D-CountDown | 0.442 | 0.419 | 0.387 | 0.555 | 0.563 | 0.520 | 0.526 | 0.457 | 0.437 | 0.499 | 0.445 | 0.417 |
5 Experiments
Experimental Setup
We evaluate the proposed methods against other sparse activation baselines, primarily CATS (Lee et al., 2024) and Deja Vu (Liu et al., 2023). We also include a Dense variant without any sparse activation for comparison. Experiments are conducted using four diverse state-of-the-art LLMs ranging from 8B to 14B parameters: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), gemma-2-9b-it (Gemma Team et al., 2024), Qwen2.5-14B-Instruct (Qwen et al., 2024), and phi-4 (Abdin et al., 2024). We test multiple sparsity ratios by varying from 0.7 to 0.9, representing the fraction of parameters excluded from computation. Implementation details are provided in Appendix A.2.
We examine both model performance preservation and computational efficiency. For model performance, we use the lm-eval-harness (Gao et al., 2024) framework to assess downstream tasks including ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), OpenbookQA (Mihaylov et al., 2018), TruthfulQA (Lin et al., 2022), WinoGrande (Sakaguchi et al., 2020), and GSM8K (Cobbe et al., 2021). Unlike prior sparse activation studies, we also evaluate conversational ability using LLM-as-a-Judge framework AlpacaEval 2.0 (Li et al., 2023a).
To assess computational efficiency and inference speed, we benchmark kernel-level latency to quantify Gated-MLP speedups from sparse activation. We also measure end-to-end token throughput and analyze theoretical reductions in floating-point operations (FLOPs) and memory traffic.
Downstream Task Performance
As shown in Table 1, in the SPIdeal setting,D-CountDown consistently outperforms all methods across all models and sparsity ratios, exhibiting negligible degradation even when compared to the dense baseline. This demonstrates the effectiveness of D-CountDown’s sparsity criterion: the indicator accurately reflects each parameter’s importance to the final output, serving as the coefficient in our linear combination formulation. This provides more informed filtering than methods like CATS which rely solely on gating magnitude. Even at 90% sparsity, D-CountDown retains only the most impactful neurons, limiting performance drop to 5.5% in the best case among evaluated models.
M-CountDown, although less effective than D-CountDown, consistently outperforms CATS. The gap between the two widens as the sparsity ratio increases, reaching over 29.4%. This demonstrates that M-CountDown ’s indicator is more predictive of useful computation than CATS’ indicator . This may seem counterintuitive since and contribute symmetrically via their elementwise product and thus should be equally informative. We revisit this comparison in section 6.
Deja Vu, which assumes ReLU-style zero-out behavior, suffers severe degradation in the SPIdeal setting. Given its reliance on predictors, which would further degrade under the SPPrac setting, we excluded it from subsequent experiments.
In the SPPrac setting, D-CountDown experiences performance loss relative to the SPIdeal due to predictor sub-optimality, suggesting better prediction strategies are needed to fully realize its potential in deployment. In contrast, M-CountDown, thanks to its predictor-free design, exhibits nearly identical performance to its SPIdeal counterpart. Notably, M-CountDown continues to outperform CATS across all sparsity settings, reinforcing the effectiveness of its signal even under realistic constraints.
LLM Chat Performance
While prior studies rely on downstream task accuracy or perplexity, these metrics often fail to capture conversational performance. To address this, we evaluate each method using an LLM-as-a-Judge framework that directly assesses chat-level performance.
As shown in Table 2, M-CountDown maintains nearly identical performance between the SPIdeal and SPPrac settings, while also outperforming CATS in both. D-CountDown exhibits noticeable degradation in SPPrac due to predictor limitations, but retains a dominant lead under SPIdeal. This trend aligns with the results observed in the downstream task evaluations.
InferenceMode | AlpacaEval 2.0 | ||
SPIdeal | |||
CATS | 25.10 | 1.72 | 0.19 |
M-CountDown | 45.84 | 29.22 | 3.90 |
D-CountDown | 48.86 | 45.85 | 29.95 |
SPPrac | |||
CATS | 31.63 | 10.47 | 0.25 |
M-CountDown | 38.31 | 33.80 | 15.88 |
D-CountDown | 3.40 | 2.81 | 1.16 |
Efficiency and Speed
To confirm that reductions in computation indeed translate into inference speedups, we measured kernel-level execution latency under various sparsity ratios. Each kernel’s execution time was recorded from the start of the Gated-MLP computation, explicitly excluding other operations like token embedding or attention mechanisms. This allowed us to isolate the precise efficiency gains attributable to sparse activation.
As shown in Figure 3, D-CountDown achieves the fastest kernel execution time overall, despite the presence of a predictor, by skipping all three weight matrix computations. Although both M-CountDown and CATS are predictor-free, M-CountDown slightly outperforms CATS in kernel speed. Given that the only architectural difference between their kernels is whether the non-linear activation function is fused, this suggests that M-CountDown gains a minor but consistent speed advantage by fusing the activation computation, thereby reducing memory traffic and avoiding additional overhead.
Furthermore, we measured average tokens generated per second for generation lengths of 512 and 1024, providing a model-level speedup assessment in typical generation scenarios. As shown in Table 3, M-CountDown achieves the highest end-to-end token throughput. Meanwhile, D-CountDown demonstrates the best performance at the kernel level, and with further optimization, its overall throughput may be further enhanced.
6 Analysis
M-CountDown vs CATS
While CATS and M-CountDown share similar core ideas for sparse activation, our experimental results show that M-CountDown consistently achieves better performance. To understand the performance gap between the indirect coefficient vectors and , we conduct a comparative analysis of how each influences and aligns with the oracle-like reference signal , the direct coefficient used in D-CountDown.
To enable direct comparison, we define binary masks , , and based on the top- magnitude entries of each vector. Each mask marks components as “alive” (1) if they survive quantile thresholding, and "dead" (0) otherwise. These binary masks are equivalent to the index sets used for sparse activation, as each represents the support of the corresponding in vector form.
Method | FLOPs(M) | Mem.(MB) | Throughput | ||
0.0 | Dense | 352.41 | 168.121 | 24.64 | 22.63 |
0.7 | CATS | 188.00 | 89.746 | 32.62 | 29.40 |
MC | 187.95 | 89.719 | 33.61 | 30.32 | |
DC | 124.59 | 59.480 | 30.69 | 27.80 | |
0.8 | CATS | 164.52 | 78.550 | 32.72 | 29.60 |
MC | 164.46 | 78.522 | 33.80 | 30.61 | |
DC | 89.37 | 42.684 | 30.70 | 27.57 | |
0.9 | CATS | 141.02 | 67.345 | 32.98 | 29.81 |
MC | 140.96 | 67.318 | 33.51 | 30.78 | |
DC | 54.11 | 25.877 | 30.73 | 27.55 |
We first define a metric called Comparative Influential Factor (CIF) to measure how much influence (or ) has on the final decision of , especially in cases where it overrides the other component. Analogously, for instance, measures how often “rescues” a component that would otherwise have been pruned by , allowing it to survive in due to its strong contribution. Formally, this is computed as:
This formulation follows from the definition of as the elementwise product of and . When is alive but is small enough to be pruned, it implies that must have been large enough to compensate, effectively “saving” that entry.
Next, we define the Comparative Agreement Factor (CAF) to evaluate how often one signal aligns with while the other disagrees. For instance, measures how frequently agrees with on keeping a component, specifically when disagrees. This is given by:
Both CIF and CAF can also be defined symmetrically for the “dead” case by inverting the roles of activation and pruning.
As shown in Figure 4, outperforms across all sparsity levels in both CIF and CAF. These results suggest that more closely reflects the true activation behavior captured by and exerts a greater direct impact on sparsity decisions than . In other words, is more effective at preserving important activations and filtering out unimportant ones, explaining M-CountDown’s stable and reliable performance under sparsity.
Nevertheless, since M-CountDown still relies on an indirect coefficient , it cannot fully match the upper-bound performance of D-CountDown, which uses the full signal directly. Despite ’s strong CIF and CAF scores, substantial mismatches with respect to the oracle mask remain, with peak CIF values reaching only about 0.6 and CAF values about 0.4, underscoring the need for future work to translate D-CountDown’s upper-bound potential into SPPrac deployments.
Possible Predictor Candidate: TernaryLinear
D-CountDown demonstrates a theoretically sound and effective sparse activation strategy, achieving strong performance in the SPIdeal setting. However, in the SPPrac scenario, performance degradation occurs due to the predictor’s limited accuracy in recovering optimal sparsity patterns. This reflects the difficulty of the prediction task rather than a flaw in the sparsity criterion itself. The task simultaneously demands precision and computational efficiency, presenting a significant challenge with considerable room for improvement.
To empirically explore this potential, we evaluate an alternative predictor architecture, TernaryLinear, whose weights are quantized as . We compare its performance with the previously utilized low-rank approximator. TernaryLinear achieves significant parameter compression by sacrificing numerical precision while preserving the matrix rank structure. Motivated by recent studies demonstrating successful LLM pretraining with ternary quantization while retaining strong model performance (Ma et al., 2024), we regard TernaryLinear as a promising candidate due to its demonstrated expressiveness even under aggressive quantization.
As shown in Table 4, TernaryLinear outperforms the low-rank baseline in F1 score, while also being more compact in terms of memory footprint. This suggests that preserving rank information, even at the cost of numerical precision, is more effective for sparse mask recovery than the reverse approach.
However, TernaryLinear has not yet been adopted due to its relatively slower runtime despite its small size. This limitation stems not from algorithmic complexity, but rather the lack of optimized GPU kernel support for ultra-low-precision operations. Prior work (Ma et al., 2025) suggests that future advances in kernel optimization and ultra-low-bit quantization are needed to fully leverage such architectures. With these improvements, techniques like TernaryLinear could become viable candidates for enabling D-CountDown to achieve its full SPIdeal performance in SPPrac scenarios.
Metric | TernaryLinear | Low-Rank |
Latency (ms) | 0.082 | 0.030 |
Theoretical footprint (MiB) | 112 | 144 |
F1-score (%) | 0.435 | 0.403 |
7 Conclusion
We introduce CountDown, a novel sparse activation framework for improving inference efficiency of large language models. To overcome the limitations of traditional non-linear activation-based sparsity, we reformulate the computation as a weighted sum over the FFNN’s down projection matrix, effectively capturing inherent sparsity in modern Gated-MLP architectures. From this perspective, we present two complementary strategies: M-CountDown, which uses derived from a single matrix as its activation indicator, achieves faster inference and better performance preservation than prior state-of-the-art methods while remaining predictor-free. D-CountDown directly leverages , the coefficient vector of the weighted sum, for fine-grained sparsity selection, demonstrating robust performance even when skipping 90% of computations under ideal conditions.
Limitations
Like most prior work on sparse activation, our study assumes a single-batch greedy decoding setting in on-device environments. While this scenario is realistic for latency-sensitive edge inference, it may be less applicable in multi-batch or server-based deployments. In such cases, strategies such as computing the union of predicted index sets across multiple samples could be explored. However, such an approach would require further investigation into how much parameter activation can be shared across inputs, a direction we leave for future work.
Additionally, our sparsity criteria rely exclusively on activation magnitude. This choice offers clear interpretability and aligns well with the weighted-sum perspective we adopt. Nevertheless, alternative sparsity metrics, such as those explored by (Akhauri et al., 2024), remain an open research avenue. Expanding beyond simple magnitude-based thresholding could further enhance the performance of sparse activation methods.
Ethical Considerations
We affirm adherence to the ACL Rolling Review (ARR) ethical guidelines, explicitly addressing potential risks and responsible research practices. This research focuses on optimizing computational efficiency in large language models (LLMs), aimed at reducing resource usage and consequently lowering environmental impact. We foresee no direct risks or potential harms to individuals or communities resulting from this work.
Comprehensive details regarding the ethical use of scientific artifacts, reproducibility of computational experiments, and related considerations are thoroughly documented in Appendix A.
Appendix A Experimental Details
A.1 Hyperparameters
Name Low-Rank, BitLinear Loss Binary Cross Entropy Sparsity ratio () Learning rate {1e-3, 5e-4} Training batch size Optimizer AdamW Target Binary mask Epochs Seed Predictor shape Low-Rank Approximator TernaryLinear – Hardware 1 NVIDIA A100 80GB
A.2 Environments
All experiments were performed on an NVIDIA A100 80GB GPU. We used Triton v3.1.0 for custom kernel development, while the rest of the experimental pipeline was built on HuggingFace Transformers v4.51.3, PyTorch v2.5.1, and CUDA v12.1.
A.3 Dataset Description
Table 6 summarizes the licenses and dataset statistics used for evaluation.
We evaluate seven Natural Language Understanding(NLU) tasks and one Natural Language Generation(NLG) task focused on mathematical reasoning (GSM8K). All datasets primarily contain English text.
Dataset | License | Train | Test |
ARC-Easy | cc-by-sa-4.0 | 2251 (500) | 2376 |
ARC-Challenge | cc-by-sa-4.0 | 1119 (500) | 1172 |
HellaSwag | MIT | 39905 (500) | 10042 |
PIQA | AFL 3.0 | 16113 (500) | 1838 |
WinoGrande | apache-2.0 | 40398 (500) | 1267 |
OpenBookQA | apache-2.0 | 4957 (500) | 500 |
TruthfulQA | apache-2.0 | 0 | 817 |
GSM8K | MIT | 0 | 1319 |
Appendix B Pseudo Codes
B.1 Procedures for Training Predictor
B.2 Kernel in Detail: M-CountDown
B.3 Kernel in Detail: D-CountDown
Appendix C Full Results
All downstream task results are in Table 11 and Table 12. Chat performance results are shown in Table 7. Kernel Speed results are shown in Figure 5. SparsityReal indicates the actual proportion of indicator elements filtered out during SPPrac inferences.
Method | Scenario | Llama-3.1-8B | Gemma-2-9B | ||||
Target Sparsity | Target Sparsity | ||||||
0.70 | 0.80 | 0.90 | 0.70 | 0.80 | 0.90 | ||
CATS | SPIdeal | 1.02 | 0.48 | 0.50 | 35.41 | 2.55 | 0.00 |
SPPrac (Win) | 3.26 | 0.55 | 0.72 | 40.76 | 6.72 | 0.00 | |
SPPrac (SparsityReal) | 70.8 | 80.0 | 89.7 | 68.8 | 79.3 | 88.1 | |
DC | SPIdeal | 45.79 | 39.33 | 11.85 | 50.44 | 48.90 | 37.79 |
SPPrac (Win) | 1.35 | 1.57 | 0.77 | 6.72 | 7.99 | 3.57 | |
SPPrac (SparsityReal) | 68.8 | 71.5 | 80.8 | 67.0 | 72.9 | 83.2 | |
MC | SPIdeal | 46.59 | 2.74 | 0.60 | 47.81 | 41.83 | 6.91 |
SPPrac (Win) | 9.68 | 3.19 | 0.74 | 48.78 | 47.88 | 28.08 | |
SPPrac (SparsityReal) | 72.7 | 82.3 | 91.0 | 68.0 | 77.8 | 87.6 |
Method | Scenario | Qwen2.5-14B | Phi-4 | ||||
Target Sparsity | Target Sparsity | ||||||
0.70 | 0.80 | 0.90 | 0.70 | 0.80 | 0.90 | ||
CATS | SPIdeal | 21.94 | 0.38 | 0.00 | 42.05 | 3.45 | 0.25 |
SPPrac (Win) | 33.62 | 7.80 | 0.00 | 48.87 | 26.81 | 0.31 | |
SPPrac (SparsityReal) | 70.0 | 80.0 | 89.1 | 67.6 | 78.3 | 89.7 | |
DC | SPIdeal | 50.10 | 48.71 | 32.73 | 49.10 | 46.46 | 37.45 |
SPPrac (Win) | 4.40 | 0.77 | 0.20 | 1.11 | 0.90 | 0.12 | |
SPPrac (SparsityReal) | 66.6 | 82.8 | 87.6 | 65.4 | 78.5 | 86.8 | |
MC | SPIdeal | 45.01 | 36.90 | 2.83 | 43.96 | 35.43 | 5.28 |
SPPrac (Win) | 48.57 | 42.51 | 15.89 | 46.24 | 41.62 | 18.82 | |
SPPrac (SparsityReal) | 70.0 | 80.0 | 90.0 | 70.1 | 79.6 | 89.6 |
Appendix D Theoretical Analysis Details
D.1 Notation
Notation | Explanation |
act FLOPs (e.g. SiLU ) |
D.2 heoretical FLOPs Analysis
Method | Compute | Explanation |
Dense | ||
CATS | ||
M-CountDown | ||
D-CountDown |
D.3 Theoretical Memory Traffic Analysis
Method | Mem. R/W | Explanation |
Dense | ||
CATS |
Method | Mem. R/W | Explanation |
M-Countdown | ||
D-Countdown |
Sparsity | Method | ARC-C | TFQA | HS | ARC-E | PIQA | WG | OBQA | GSM8K |
Llama-3.1-8B-Instruct | |||||||||
0.0 | DENSE | 0.520 | 0.367 | 0.590 | 0.819 | 0.800 | 0.737 | 0.336 | 0.760 |
0.7 | DEJAVU | 0.292 | 0.229 | 0.272 | 0.445 | 0.553 | 0.503 | 0.218 | 0.000 |
CATS | 0.453 | 0.343 | 0.523 | 0.754 | 0.739 | 0.653 | 0.298 | 0.003 | |
M-COUNTDOWN | 0.493 | 0.372 | 0.568 | 0.784 | 0.776 | 0.695 | 0.330 | 0.544 | |
D-COUNTDOWN | 0.509 | 0.370 | 0.592 | 0.812 | 0.795 | 0.727 | 0.332 | 0.688 | |
0.8 | DEJAVU | 0.282 | 0.231 | 0.273 | 0.440 | 0.557 | 0.511 | 0.228 | 0.000 |
CATS | 0.358 | 0.326 | 0.428 | 0.651 | 0.676 | 0.582 | 0.278 | 0.000 | |
M-COUNTDOWN | 0.458 | 0.343 | 0.534 | 0.759 | 0.748 | 0.661 | 0.314 | 0.288 | |
D-COUNTDOWN | 0.502 | 0.356 | 0.585 | 0.809 | 0.789 | 0.713 | 0.334 | 0.605 | |
0.9 | DEJAVU | 0.296 | 0.230 | 0.273 | 0.455 | 0.557 | 0.530 | 0.236 | 0.000 |
CATS | 0.293 | 0.252 | 0.303 | 0.495 | 0.574 | 0.534 | 0.242 | 0.000 | |
M-COUNTDOWN | 0.411 | 0.304 | 0.430 | 0.649 | 0.676 | 0.613 | 0.286 | 0.001 | |
D-COUNTDOWN | 0.484 | 0.330 | 0.548 | 0.776 | 0.755 | 0.680 | 0.312 | 0.313 | |
Qwen2.5-14B-Instruct | |||||||||
0.0 | DENSE | 0.608 | 0.517 | 0.657 | 0.861 | 0.817 | 0.758 | 0.364 | 0.807 |
0.7 | DEJAVU | 0.336 | 0.318 | 0.365 | 0.612 | 0.616 | 0.533 | 0.254 | 0.000 |
CATS | 0.488 | 0.443 | 0.585 | 0.777 | 0.729 | 0.629 | 0.318 | 0.043 | |
M-COUNTDOWN | 0.573 | 0.488 | 0.638 | 0.829 | 0.792 | 0.704 | 0.352 | 0.776 | |
D-COUNTDOWN | 0.588 | 0.518 | 0.654 | 0.850 | 0.801 | 0.736 | 0.364 | 0.770 | |
0.8 | DEJAVU | 0.340 | 0.322 | 0.357 | 0.609 | 0.619 | 0.554 | 0.258 | 0.000 |
CATS | 0.410 | 0.371 | 0.472 | 0.683 | 0.632 | 0.568 | 0.284 | 0.000 | |
M-COUNTDOWN | 0.532 | 0.476 | 0.614 | 0.813 | 0.743 | 0.670 | 0.352 | 0.681 | |
D-COUNTDOWN | 0.579 | 0.488 | 0.644 | 0.837 | 0.799 | 0.716 | 0.360 | 0.751 | |
0.9 | DEJAVU | 0.358 | 0.333 | 0.369 | 0.612 | 0.621 | 0.531 | 0.256 | 0.000 |
CATS | 0.356 | 0.327 | 0.385 | 0.619 | 0.621 | 0.547 | 0.260 | 0.000 | |
M-COUNTDOWN | 0.468 | 0.421 | 0.525 | 0.736 | 0.686 | 0.589 | 0.304 | 0.100 | |
D-COUNTDOWN | 0.512 | 0.436 | 0.607 | 0.801 | 0.756 | 0.648 | 0.312 | 0.371 | |
gemma-2-9b-it | |||||||||
0.0 | DENSE | 0.632 | 0.433 | 0.597 | 0.856 | 0.812 | 0.761 | 0.404 | 0.663 |
0.7 | DEJAVU | 0.339 | 0.246 | 0.300 | 0.596 | 0.590 | 0.532 | 0.276 | 0.000 |
CATS | 0.575 | 0.412 | 0.559 | 0.840 | 0.755 | 0.680 | 0.348 | 0.565 | |
M-COUNTDOWN | 0.605 | 0.421 | 0.592 | 0.849 | 0.793 | 0.726 | 0.374 | 0.632 | |
D-COUNTDOWN | 0.626 | 0.417 | 0.600 | 0.854 | 0.800 | 0.750 | 0.384 | 0.649 | |
0.8 | DEJAVU | 0.346 | 0.246 | 0.296 | 0.599 | 0.581 | 0.548 | 0.262 | 0.000 |
CATS | 0.490 | 0.366 | 0.486 | 0.788 | 0.696 | 0.604 | 0.328 | 0.105 | |
M-COUNTDOWN | 0.583 | 0.408 | 0.582 | 0.842 | 0.767 | 0.707 | 0.360 | 0.610 | |
D-COUNTDOWN | 0.604 | 0.421 | 0.599 | 0.851 | 0.796 | 0.728 | 0.374 | 0.624 | |
0.9 | DEJAVU | 0.356 | 0.246 | 0.303 | 0.616 | 0.573 | 0.523 | 0.264 | 0.000 |
CATS | 0.364 | 0.242 | 0.310 | 0.617 | 0.589 | 0.537 | 0.278 | 0.000 | |
M-COUNTDOWN | 0.534 | 0.383 | 0.517 | 0.799 | 0.727 | 0.648 | 0.344 | 0.438 | |
D-COUNTDOWN | 0.578 | 0.410 | 0.572 | 0.833 | 0.777 | 0.676 | 0.352 | 0.524 | |
phi-4 | |||||||||
0.0 | DENSE | 0.558 | 0.404 | 0.632 | 0.814 | 0.808 | 0.766 | 0.338 | 0.923 |
0.7 | DEJAVU | 0.387 | 0.311 | 0.348 | 0.655 | 0.626 | 0.587 | 0.266 | 0.000 |
CATS | 0.536 | 0.400 | 0.595 | 0.794 | 0.791 | 0.696 | 0.304 | 0.807 | |
M-COUNTDOWN | 0.533 | 0.384 | 0.616 | 0.800 | 0.796 | 0.733 | 0.334 | 0.888 | |
D-COUNTDOWN | 0.554 | 0.411 | 0.630 | 0.809 | 0.807 | 0.752 | 0.332 | 0.916 | |
0.8 | DEJAVU | 0.409 | 0.333 | 0.354 | 0.655 | 0.632 | 0.585 | 0.270 | 0.000 |
CATS | 0.516 | 0.397 | 0.539 | 0.771 | 0.760 | 0.644 | 0.298 | 0.351 | |
M-COUNTDOWN | 0.503 | 0.386 | 0.594 | 0.792 | 0.778 | 0.715 | 0.330 | 0.767 | |
D-COUNTDOWN | 0.552 | 0.408 | 0.622 | 0.807 | 0.810 | 0.755 | 0.340 | 0.898 | |
0.9 | DEJAVU | 0.392 | 0.317 | 0.357 | 0.640 | 0.630 | 0.566 | 0.266 | 0.000 |
CATS | 0.426 | 0.356 | 0.414 | 0.676 | 0.672 | 0.591 | 0.280 | 0.000 | |
M-COUNTDOWN | 0.479 | 0.370 | 0.524 | 0.759 | 0.728 | 0.654 | 0.296 | 0.287 | |
D-COUNTDOWN | 0.529 | 0.399 | 0.601 | 0.798 | 0.789 | 0.695 | 0.318 | 0.827 |
Sparsity | Method | SparsityReal | ARC-C | TFQA | HS | ARC-E | PIQA | WG | OBQA | GSM8K |
Llama-3.1-8B-Instruct | ||||||||||
0.7 | CATS | 0.684 | 0.461 | 0.355 | 0.549 | 0.778 | 0.764 | 0.683 | 0.316 | 0.127 |
M-COUNTDOWN | 0.709 | 0.484 | 0.375 | 0.574 | 0.788 | 0.778 | 0.708 | 0.310 | 0.547 | |
D-COUNTDOWN | 0.705 | 0.422 | 0.318 | 0.373 | 0.748 | 0.714 | 0.663 | 0.298 | 0.002 | |
0.8 | CATS | 0.784 | 0.420 | 0.322 | 0.495 | 0.718 | 0.721 | 0.624 | 0.296 | 0.000 |
M-COUNTDOWN | 0.806 | 0.460 | 0.361 | 0.549 | 0.770 | 0.757 | 0.680 | 0.322 | 0.322 | |
D-COUNTDOWN | 0.739 | 0.382 | 0.306 | 0.388 | 0.688 | 0.673 | 0.621 | 0.292 | 0.003 | |
0.9 | CATS | 0.902 | 0.299 | 0.273 | 0.323 | 0.521 | 0.607 | 0.537 | 0.238 | 0.000 |
M-COUNTDOWN | 0.895 | 0.416 | 0.321 | 0.471 | 0.711 | 0.721 | 0.620 | 0.304 | 0.009 | |
D-COUNTDOWN | 0.843 | 0.349 | 0.285 | 0.345 | 0.628 | 0.635 | 0.593 | 0.260 | 0.000 | |
Qwen2.5-14B-Instruct | ||||||||||
0.7 | CATS | 0.698 | 0.518 | 0.460 | 0.612 | 0.805 | 0.761 | 0.660 | 0.336 | 0.293 |
M-COUNTDOWN | 0.719 | 0.590 | 0.509 | 0.640 | 0.838 | 0.792 | 0.712 | 0.358 | 0.767 | |
D-COUNTDOWN | 0.678 | 0.513 | 0.426 | 0.536 | 0.798 | 0.748 | 0.668 | 0.322 | 0.197 | |
0.8 | CATS | 0.802 | 0.472 | 0.421 | 0.551 | 0.754 | 0.712 | 0.627 | 0.284 | 0.000 |
M-COUNTDOWN | 0.804 | 0.553 | 0.492 | 0.625 | 0.826 | 0.769 | 0.669 | 0.354 | 0.704 | |
D-COUNTDOWN | 0.827 | 0.454 | 0.394 | 0.468 | 0.740 | 0.693 | 0.615 | 0.292 | 0.000 | |
0.9 | CATS | 0.906 | 0.347 | 0.350 | 0.393 | 0.631 | 0.616 | 0.527 | 0.258 | 0.000 |
M-COUNTDOWN | 0.889 | 0.492 | 0.450 | 0.580 | 0.794 | 0.727 | 0.632 | 0.320 | 0.287 | |
D-COUNTDOWN | 0.893 | 0.434 | 0.384 | 0.429 | 0.689 | 0.669 | 0.605 | 0.282 | 0.000 | |
gemma-2-9b-it | ||||||||||
0.7 | CATS | 0.695 | 0.580 | 0.427 | 0.567 | 0.843 | 0.770 | 0.693 | 0.368 | 0.593 |
M-COUNTDOWN | 0.685 | 0.608 | 0.431 | 0.598 | 0.854 | 0.801 | 0.745 | 0.386 | 0.633 | |
D-COUNTDOWN | 0.689 | 0.567 | 0.403 | 0.493 | 0.821 | 0.751 | 0.702 | 0.364 | 0.340 | |
0.8 | CATS | 0.806 | 0.542 | 0.392 | 0.501 | 0.811 | 0.729 | 0.615 | 0.346 | 0.083 |
M-COUNTDOWN | 0.779 | 0.596 | 0.412 | 0.589 | 0.847 | 0.788 | 0.712 | 0.370 | 0.618 | |
D-COUNTDOWN | 0.755 | 0.564 | 0.401 | 0.506 | 0.819 | 0.758 | 0.702 | 0.374 | 0.381 | |
0.9 | CATS | 0.911 | 0.340 | 0.258 | 0.306 | 0.617 | 0.586 | 0.514 | 0.262 | 0.000 |
M-COUNTDOWN | 0.875 | 0.573 | 0.395 | 0.554 | 0.829 | 0.761 | 0.686 | 0.360 | 0.544 | |
D-COUNTDOWN | 0.853 | 0.529 | 0.383 | 0.492 | 0.806 | 0.747 | 0.665 | 0.354 | 0.187 | |
phi-4 | ||||||||||
0.7 | CATS | 0.675 | 0.539 | 0.417 | 0.613 | 0.801 | 0.795 | 0.724 | 0.322 | 0.856 |
M-COUNTDOWN | 0.707 | 0.540 | 0.393 | 0.620 | 0.804 | 0.796 | 0.736 | 0.332 | 0.894 | |
D-COUNTDOWN | 0.687 | 0.471 | 0.368 | 0.485 | 0.750 | 0.733 | 0.685 | 0.294 | 0.208 | |
0.8 | CATS | 0.771 | 0.525 | 0.390 | 0.587 | 0.795 | 0.786 | 0.673 | 0.300 | 0.675 |
M-COUNTDOWN | 0.799 | 0.527 | 0.381 | 0.607 | 0.793 | 0.784 | 0.715 | 0.334 | 0.817 | |
D-COUNTDOWN | 0.815 | 0.418 | 0.343 | 0.438 | 0.707 | 0.692 | 0.657 | 0.268 | 0.036 | |
0.9 | CATS | 0.889 | 0.458 | 0.360 | 0.460 | 0.713 | 0.692 | 0.605 | 0.294 | 0.000 |
M-COUNTDOWN | 0.894 | 0.498 | 0.386 | 0.563 | 0.777 | 0.750 | 0.698 | 0.316 | 0.450 | |
D-COUNTDOWN | 0.895 | 0.408 | 0.294 | 0.404 | 0.674 | 0.668 | 0.616 | 0.270 | 0.000 |